# Efficient Dual-domain Image Dehazing with Haze Prior Perception

### v1

@@ IV-C2 Effectiveness of HAFM
According to Table IV, using only spatial (HAFM-S) or only frequency modulation (HAFM-F) leads to PSNR drops of 1.06 dB and 1.00 dB, respectively, compared to the full HAFM. This highlights the necessity of combining both spatial and frequency modulation for optimal performance. When dark channel-guided attention is replaced with standard spatial attention (HAFM-SSA), PSNR decreases by 0.40 dB, and SSIM drops by 0.001. This confirms that dark channel-guided attention provides more effective haze localization than its conventional counterpart. Furthermore, replacing the frequency-domain MLP with a spatial MLP (HAFM-SMLP) results in a 48.1% increase in FLOPs, yet PSNR remains 0.27 dB lower than the full HAFM. This demonstrates that frequency-domain processing not only improves computational efficiency but also enhances global feature modeling.
TABLE V: Ablation study for key components in MGAM.
@@
IV-C3 Effectiveness of MGAM
To assess the contribution of each MGAM component, we conduct an ablation study with four variants: MGAM-$3\times 3$ (only the $3\times 3$ gated branch), MGAM-$5\times 5$ (only the $5\times 5$ gated branch), MGAM-Nogate (both branches without gating), and MGAM-Noskip (removing the skip connection). The results are reported in Table V.
From the results, MGAM-$3\times 3$ achieves 0.21 dB higher PSNR but 0.001 lower SSIM compared to MGAM-$5\times 5$. This indicates that small-kernel convolutions are more sensitive to local texture recovery, whereas large-kernel convolutions are better at capturing global structural information. By integrating both branches, the full MGAM balances these strengths, yielding simultaneous improvements in PSNR and SSIM.
Removing the gating mechanism while retaining the dual-branch structure (MGAM-Nogate) causes PSNR and SSIM to drop by 1.03 dB and 0.001, respectively, highlighting the importance of dynamic gating. By allowing low-level features to regulate the flow of high-level information, the gating mechanism enhances adaptive feature selection and helps mitigate detail loss in deeper layers. Additionally, removing the skip connection from the original input (MGAM-Noskip) results in a PSNR decrease of 0.42 dB and an SSIM decrease of 0.001, underscoring the importance of the skip connection in correcting selection bias introduced by the gating mechanism.
@@ IV-C3 Effectiveness of MGAM
(a) Hazy image (b) GT (c) Dark channel (d) Attn map of the first HAFM (e) Attn map of the last HAFM
IV-C4 Effectiveness of PCGB
To validate the rationale behind PCGB, we implement four variants: (1) PCGB-SSA replaces dark channel-guided spatial attention with standard spatial attention, the same as HAFM-SSA; (2) PCGB-NoFR generates the spatial attention map using only the initial dark channel features, without feedback refinement; (3) PCGB-DAF directly adds feedback correction features to the initial dark channel features instead of using SKFusion [14]; (4) PCGB-PFF uses the initial prior only in the first DGFDBlock, while subsequent blocks fuse the previous guidance with correction features via SKFusion. Table VI and Table VII present the results on SOTS-Indoor and SOTS-outdoor datasets, respectively.
TABLE VI: Ablation study for key components in PCGB on SOTS-Indoor.
@@ IV-C4 Effectiveness of PCGB
TABLE VII: Ablation study for key components in PCGB on SOTS-Outdoor.
@@
Indoor results. Results in Table VI indicate that the dark channel prior is relatively reliable in indoor scenes, making it an effective guidance cue. Even without feedback refinement, PCGB-NoFR outperforms PCGB-SSA by 0.12 dB in PSNR, validating the prior’s ability to localize haze and estimate degradation. Introducing feedback correction via simple addition (PCGB-DAF) brings a further 0.12 dB gain, while replacing addition with SKFusion in the full PCGB improves PSNR and SSIM by 0.16 dB and 0.001, respectively, showing enhanced adaptability to structural complexity. However, PCGB-PFF, which discards the prior after the first block, leads to a 0.71 dB PSNR drop, suggesting that consistent use of the initial prior across layers is important for stable and accurate guidance indoors.
Outdoor results. In contrast, the dark channel prior becomes less reliable in outdoor environments due to factors such as sky and bright regions. As shown in Table VII, PCGB-NoFR performs 0.28 dB worse than PCGB-SSA, highlighting the limitations of the prior alone. Introducing feedback correction (PCGB-DAF) compensates for this, achieving gains of 0.42 dB over PCGB-NoFR and 0.14 dB over PCGB-SSA. PCGB-PFF, which progressively fuses the prior with feedback features, also improves over NoFR by 0.29 dB and performs comparably to PCGB-SSA, suggesting that reducing reliance on noisy priors is beneficial outdoors. Nevertheless, PCGB-DAF still outperforms PCGB-PFF by 0.13 dB, indicating that retaining some contribution from the initial prior remains helpful. The full PCGB achieves the best performance, surpassing PCGB-DAF by 0.56 dB, confirming that SKFusion effectively handles the diverse and non-uniform haze patterns typical of outdoor scenes.
@@ IV-C5 Visual Analysis of PCGB
In the indoor scene, the initial attention map closely reflects the raw dark channel distribution, resulting in widespread activation across the image. After feedback refinement, the attention becomes more concentrated on heavily degraded areas, improving both the accuracy of haze localization and the effectiveness of spatial modulation.
In contrast, the outdoor scene reveals the limitations of dark channel priors. Due to low dark channel values in bright areas such as the sky, the initial attention map tends to overemphasize these regions while neglecting genuinely hazy areas like distant buildings and road surfaces. Through feedback refinement, PCGB progressively redirects attention away from the sky and toward the actual degradation, narrowing the response gap and mitigating over-processing. This leads to more balanced and accurate spatial attention.
Under complex real-world degradations, PCGB exhibits even stronger corrective capabilities. In the NH-HAZE example, the initial attention map mistakenly activates on irrelevant structures such as the swing’s mesh, resulting in false positives. Moreover, it primarily focuses on dense haze areas while neglecting mildly degraded regions, leading to insufficient coverage. After iterative refinement, the final attention map more accurately aligns with the actual haze distribution. It effectively suppresses false activations and enhances sensitivity to previously neglected regions. These results confirm the robustness of PCGB in handling spatially diverse and challenging degradation patterns.
@@ V Conclusion
In this paper, we propose DGFDNet, a dual-domain dehazing framework that explicitly aligns spatial and frequency-domain degradation cues under dark channel guidance. To address the inefficiency and weak coupling in existing spatial-frequency models, DGFDNet introduces three key modules: the Haze-Aware Frequency Modulator (HAFM) for degradation-aware spectral filtering, the Multi-level Gating Aggregation Module (MGAM) for adaptive multi-scale spatial feature fusion, and the Prior Correction Guidance Branch (PCGB) for iterative refinement of dark channel priors. These modules form a compact and computationally efficient framework that effectively captures long-range dependencies while preserving fine details, robustly handling both homogeneous and non-homogeneous haze distributions. Extensive experiments on both synthetic and real-world dehazing benchmarks demonstrate that DGFDNet achieves superior dehazing performance with strong generalization and computational efficiency, making it well-suited for practical deployment.

### v2

@@ IV-C Ablation Studies
TABLE IV: Ablation study for key components in MGAM.
@@
IV-C1 Effectiveness of Core Modules
To thoroughly evaluate the effectiveness and complementarity of the three core modules, we design four variants: MGAM, HAFM (without feedback correction), HAFM+PCGB (with full feedback-guided correction), and MGAM+HAFM. Table II summarizes the performance of each variant.
As shown in Table II, MGAM improves PSNR by 0.75 dB over HAFM while maintaining the same SSIM of 0.994, highlighting its advantage in enhancing local details. Introducing PCGB into HAFM yields an additional 0.70 dB gain, emphasizing its role in enhancing global degradation modeling via adaptive prior correction. Combining MGAM and HAFM results in a 1.80 dB PSNR improvement over MGAM alone and 2.55 dB over HAFM, benefiting from the synergy between multi-scale spatial feature extraction and dual-domain modulation. Further adding PCGB achieves the best overall performance, confirming that the three modules complement each other and jointly contribute to both local and global restoration with minimal additional computational cost.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
(a) Hazy image (b) Dehamer [34] (c) Ours (d) GT
Figure 6: Visual comparisons on real-world hazy images from the Dense-Haze dataset. Key regions marked with red boxes are enlarged in left-to-right order and arranged vertically on the right.
@@ IV-C1 Effectiveness of Core Modules
(a) Hazy image (b) Dehamer [34] (c) FocalNet [38] (d) Ours (e) GT
IV-C2 Effectiveness of HAFM
To assess the role of HAFM, we design four variants: HAFM-S (spatial modulation only), HAFM-F (frequency modulation only), HAFM-SSA (standard spatial attention replacing dark channel-guided attention), and HAFM-SMLP (spatial MLP replacing frequency modulation). The results are presented in Table III.
According to Table III, using only spatial (HAFM-S) or only frequency modulation (HAFM-F) results in PSNR drops of 1.06 dB and 1.00 dB, respectively, compared to the full HAFM. This underscores the importance of combining both spatial and frequency modulation for optimal performance. Replacing dark channel-guided attention with standard spatial attention (HAFM-SSA) leads to a 0.40 dB drop in PSNR and a 0.001 decrease in SSIM, highlighting that the haze-specific spatial modulation offers a stronger prior to frequency modulation. Substituting the frequency-domain MLP with a spatial MLP (HAFM-SMLP) increases FLOPs by 48.1%, yet PSNR remains 0.27 dB lower than the full HAFM, showing that frequency-domain processing not only boosts computational efficiency but also enhances global feature modeling.
IV-C3 Effectiveness of MGAM
To assess the contribution of each MGAM component, we conduct an ablation study with four variants: MGAM-$3\times 3$ (only the $3\times 3$ gated branch), MGAM-$5\times 5$ (only the $5\times 5$ gated branch), MGAM-Nogate (both branches without gating), and MGAM-Noskip (removing the skip connection). The results are reported in Table IV.
TABLE V: Ablation study for key components in PCGB on SOTS-Indoor.
@@ IV-C3 Effectiveness of MGAM
TABLE VI: Ablation study for key components in PCGB on SOTS-Outdoor.
@@
MGAM-$3\times 3$ achieves 0.21 dB higher PSNR but 0.001 lower SSIM compared to MGAM-$5\times 5$, indicating that small-kernel convolutions are better at capturing local textures, while large-kernel convolutions capture global structures. The full MGAM, combining both, balances these strengths, improving both PSNR and SSIM. Removing the gating mechanism (MGAM-Nogate) results in a 1.03 dB drop in PSNR and 0.001 in SSIM, emphasizing the importance of dynamic gating for adaptive feature selection and mitigating detail loss. Finally, removing the skip connection (MGAM-Noskip) reduces PSNR by 0.42 dB and SSIM by 0.001, underscoring its role in correcting selection bias from the gating mechanism.
IV-C4 Effectiveness of PCGB
To validate the rationale behind PCGB, we test four variants: (1) PCGB-SSA, which employs standard spatial attention, identical to HAFM-SSA; (2) PCGB-NoFR, where the spatial attention map is generated from the initial dark channel features without feedback refinement (3) PCGB-DAF, which directly adds feedback correction to the initial dark channel features, bypassing SKFusion [14]; (4) PCGB-PFF, which uses the initial prior only in the first DGFDBlock, while subsequent blocks fuse the guidance and feedback from the previous DGFDBlock via SKFusion. Results on the SOTS-Indoor and SOTS-Outdoor datasets are exhibited in Table V and Table VI, respectively.
(a) Hazy image (b) GT (c) Dark channel (d) First HAFM Attnmap (e) Last HAFM Attnmap
Indoor results. Table V shows that the dark channel prior remains a reliable guidance cue in indoor scenes. Even without feedback refinement, PCGB-NoFR outperforms PCGB-SSA by 0.12 dB in PSNR, demonstrating the prior’s effectiveness in haze localization and degradation estimation. Adding feedback correction (PCGB-DAF) improves PSNR by 0.12 dB, and using SKFusion in the full PCGB increases both PSNR by 0.16 dB and SSIM by 0.001, demonstrating better adaptability to complex structures. However, PCGB-PFF, which discards the prior after the first block, results in a 0.71 dB PSNR drop, highlighting the importance of consistently using the initial prior for stable and accurate guidance.
Outdoor results. In outdoor environments, the dark channel prior becomes less reliable due to factors like sky and bright regions. As shown in Table VI, PCGB-NoFR performs 0.28 dB worse than PCGB-SSA, highlighting the limitations of the prior alone. Introducing feedback correction (PCGB-DAF) improves performance, yielding gains of 0.42 dB over PCGB-NoFR and 0.14 dB over PCGB-SSA. PCGB-PFF, which progressively fuses the prior with feedback, improves over NoFR by 0.29 dB and matches PCGB-SSA, proving the benefit of self-correction on noisy priors in outdoor scenes. However, PCGB-DAF still outperforms PCGB-PFF by 0.13 dB, emphasizing the importance of retaining the initial prior to avoid accumulating errors. The full PCGB achieves the best, surpassing PCGB-DAF by 0.56 dB, validating SKFusion’s ability to handle complex outdoor haze patterns.
@@ IV-C5 Visual Analysis of PCGB
In the indoor scene, the initial attention map follows the raw dark channel distribution, causing widespread activation. After feedback refinement, attention is focused on heavily degraded areas, improving haze localization and spatial modulation. In the outdoor scene, the limitations of dark channel priors are clear. In bright regions like the sky, the initial map overemphasizes these areas, neglecting distant buildings and roads. Feedback refinement shifts attention to the other haze regions, reducing over-processing of the sky.
In the NH-HAZE example, the initial attention map incorrectly activates on irrelevant structures, like the swing’s mesh, causing false positives. It also focuses on dense haze areas while missing subtly degraded regions. After refinement, the final map better aligns with the haze distribution, suppressing false activations and improving sensitivity to neglected regions. These results demonstrate the robustness of PCGB in handling complex, spatially diverse degradation patterns.
@@ V Conclusion
In this paper, we propose DGFDNet, a dual-domain dehazing framework that explicitly aligns spatial and frequency-domain degradation cues under dark channel guidance. To address inefficiencies and weak coupling in existing spatial-frequency models, DGFDNet introduces three key modules: the Haze-Aware Frequency Modulator (HAFM) for degradation-aware spectral filtering, the Multi-level Gating Aggregation Module (MGAM) for adaptive multi-scale spatial fusion, and the Prior Correction Guidance Branch (PCGB) for iterative dark channel prior refinement. These modules form a compact, efficient framework that captures long-range dependencies and preserves fine details, handling both homogeneous and non-homogeneous haze. Extensive experiments on synthetic and real-world benchmarks show DGFDNet’s superior performance, strong generalization, and computational efficiency, making it well-suited for practical deployment.